OpenStudy (anonymous):

three construction workers can throw bricks vertically up with speed of 4m/s , 6m/s and 12m/s respectively. working together, they can stack them on a platform of maximum height of????

OpenStudy (anonymous):

Umm the highest platform would be of the distance from the worker who throws the least velocity. So use the lowest velocity and determine the distance the brick can travel with that velocity. That is the height of the platform.

OpenStudy (anonymous):

@AravindG That's what I said. but you said it better =)

OpenStudy (aravindg):

"Umm the highest platform would be of the distance from the worker who throws the least velocity to the the highrst point " at romero a better way

OpenStudy (aravindg):

*highest

OpenStudy (vincent-lyon.fr):

I think there is a trick in that problem.

OpenStudy (callisto):

I'm thinking using energy approach for this question... KE loss = PE gain (1/2)mv^2 = mgh h = v^2 / (2g) = v^2 / 2(9.81) Put v into the equation. When v = 2, h = 4^2 / 19.62 =? This is the maximum height for the block being thrown with lowest initial velocity